# Semantics-Aligned, Curriculum-Driven, and Reasoning-Enhanced Vulnerability Repair Framework

### v1

@@ 3. Approach
Framework. These principles are realized through a two-stage training pipeline, as shown in Figure 4. In the first stage, we bootstrap the model’s reasoning ability. We employ a commercial teacher LLM to generate high-quality (reasoning, patch) examples, and then perform Supervised Fine-Tuning (SFT) on our open-source student model using a subset of this curated data selected through rejection sampling. This stage provides a strong initial LLM checkpoint to master the Reason-then-Edit paradigm. In the second stage, we refine the model’s patching ability using RL. The model is optimized through a syntactic- and semantic-aware RL optimization, which uses rewards measuring both syntactic and semantic correctness, and guided by difficulty-based curriculum learning, which gradually increases task complexity to multi-hunk fixes. This stage encourages exploration and the generation of functionally correct patches. The final, trained SeCuRepair model takes a vulnerable function and produces a structured output containing both the repair strategy and the corresponding patch code.
@@
Figure 4. Overview of the SeCuRepair training pipeline. (1) Stage 1: Reasoning-Enhanced SFT. A teacher LLM generates (reasoning, patch) candidates, which are filtered via rejection sampling to create a high-quality dataset. The student model is then fine-tuned to learn the reason-then-edit format. (2) Stage 2: Semantics-Guided RL. The SFT model is refined using RL. The policy is rewarded based on a composite score of syntactic (BLEU) and semantic (AST, DFG) similarity to the oracle patch. Stage 2 is organized with a curriculum, progressively training the model on repairs with an increasing number of hunks to master multi-hunk fixes.
3.1. Reasoning-transferred SFT
The first stage of our pipeline aims to initialize the model to follow a reason-then-edit paradigm. We posit that generating an explicit repair plan before generating patches improves both controllability and generalization. A plan forces the AVR model to first articulate the “what, where, and how” of a fix, providing a coherent guide for the subsequent patch generation. This abstract rationale guides the model to capture semantic repair patterns rather than the lexical details of a single patch.
3.1.1. Knowledge Distillation
To achieve this, the first step is to construct a dataset of (reasoning, patch) pairs by distilling reasoning knowledge from a powerful commercial teacher LLM (i.e., GPT-5 mini (gpt5mini, )). This process transfers the teacher’s advanced reasoning capabilities to our open-source student model, enabling it to infer nuanced, generalizable reasoning patterns from the generated examples. For each vulnerability in the training set, we prompt the teacher to produce one candidate (reasoning, patch) pairs. The student model (SeCuRepair) is then trained on a quality-filtered subset using a standard SFT objective. While SFT can lead to overfitting on lexical patterns (as discussed in Section 2.2), we intentionally leverage this tendency here. The model’s inclination for token-level mimicry is highly effective for learning the rigid output schema (i.e., …¡reason¿…¡/reason¿¡patch¿…¡/patch¿). The potential for overfitting on the patch content is a secondary concern at this stage, as it will be mitigated by the subsequent RL training. Our ablation studies (Section 5.3.2) also confirm that initializing with this SFT stage leads to more stable and effective RL training compared to starting from scratch.
3.1.2. Rejection Sampling
Outputs from a teacher LLM can be noisy, with flawed rationales and incorrect patches (prior work (favor, ) reports that even GPT-4o often produces low-quality patches). To ensure the student model learns only from high-quality examples, we adopt a strict rejection-sampling strategy to denoise the distilled data. Specifically, we apply a two-step filtering process to the $K$ candidates generated for each vulnerability:
  • Syntactic Filtering: We first discard any responses that violate the output schema, such as those with missing, misordered, or malformed ¡reason¿ or ¡patch¿ tags.
  • Semantic Filtering: For the remaining candidates, we adopt a strong heuristic proxy for reasoning quality: we retain a (reasoning, patch) pair only if its patch is similar to the ground-truth oracle patch with CodeBLEU ¿ 0.5. Our assumption is that the similarity to the oracle solution serves as a strong, albeit imperfect, proxy of the quality of the associated reasoning.
After rigorous filtering, we retain 484 high-quality examples distilled from the teacher LLM. This denoised dataset is then used to fine-tune SeCuRepair, yielding a high-fidelity starting policy for the subsequent reinforcement learning stage.
3.1.3. Supervised-fine-tuning Objective
After rejection sampling, we obtain a high-precision keep set $\mathcal{D}_{\text{keep}}=\{(X_{i},L_{i},r_{i},y_{i})\}$, where $r_{i}$ is the teacher’s reasoning trace. Each $r_{i}$ instance is serialized in the reason-then-edit schema: <reason> $r_{i}$ </reason> <patch> $y_{i}$ </patch>.
Following existing approaches (favor, ; vulmaster, ; vrepair, ), we fine-tune the student model with a standard SFT objective over the keep set $\mathcal{D}_{\text{keep}}$:
$\mathcal{L}_{\text{SFT}}\;=\;\frac{1}{\sum_{i}T_{i}^{R}}\sum_{i}\;\sum_{t=1}^{T_{i}^{R}}-\log\pi_{\theta}\!\big(r_{i,t},y_{i,t}\;\big|\;P_{i},\;r_{i,<t}\big),$
where $P_{i}$ is the input prompt (vulnerable function and instructions), the $R_{i}$ is the model response that includes reasoning trace and patch $(r_{i},y_{i})$, and $\theta$ represents the model’s parameters. Note that the loss is computed only on response tokens. $T_{i}^{R}$} is the total number of tokens in the response.
@@ 3.1.3. Supervised-fine-tuning Objective
The resulting model, denoted as $\theta_{\text{SFT}}$, is now capable of generating structured reasoning before patching and serves as the initial policy for the subsequent reinforcement learning stage.
3.2. Reinforcement Learning
While the SFT stage provides a format-faithful starting point for reasoning, it falls short in enabling the model to learn generalizable repairs. By training the model to mimic a single reference patch with token-level maximum likelihood, SFT conflates essential repair logic with spurious lexical styles. To overcome this, our second stage uses RL with semantics-aware rewards to refine the model. Our RL design addresses this objective mismatch by enabling the policy model, initialized from $\theta_{\text{SFT}}$, to explore the vast space of candidate patches and to learn their structural and semantic repair logic, rather than merely producing textually identical outputs.
@@ 3.2. Reinforcement Learning
We employ an on-policy RL algorithm GRPO (shao2024deepseekmath, ) for this refinement process. During training, the model generates a (reasoning, patch) pair for a given vulnerability. We then compute a scalar reward based on the generated patch, $\hat{y}_{i}$. This semantics-aware reward represents a critical component of our RL design: it moves beyond simple token overlap to measure the syntactic (Abstract Syntax Tree, AST) and semantic (Data Flow Graph, DFG) similarities between the generated patch and the oracle solution using tree-matching. This reward signal is then used to update the model’s policy, encouraging it to generate patches that achieve higher semantic agreement with the oracle. The final, tuned policy model, $\pi_{\theta}$, serves as the backend model for SeCuRepair.
3.2.1. Generation
The reinforcement learning process begins with on-policy generation. At each iteration, for each input prompt $P_{i}=(L_{i},X_{i})$ in a training batch, the current policy $\pi_{\theta}$ (initialized from $\theta_{\text{SFT}}$) generates $M$ complete (reasoning, patch) sequences. We then parse these M responses to automatically extract the set of candidate patches (i.e., $\hat{y}_{i1},\ldots,\hat{y}_{iM}$) by matching the patch tags. This collection of generated patches represents the policy’s current behavior and is then passed to our reward function for evaluation, providing the necessary signal for the policy update.
3.2.2. Reward
To guide the policy toward learning syntactic and semantic patterns for vulnerability repair, we designed a semantics-aware reward function, $Re$. This function computes a scalar score that estimates the quality of a generated patch, $\hat{y}_{i}$, by measuring its syntactic and semantic agreement with the oracle patch, $y_{i}$. A higher reward signal indicates that the generated patch is more functionally aligned with the ground-truth solution. The reward is then converted into normalized advantages and used to guide the policy update.
Specifically, our reward function assesses the agreement between the generated patch and the oracle patch at three distinct levels: 1) lexical agreement; 2) syntactic agreement; and 3) semantic agreement. The final reward is a combination of these three scores. Crucially, our reward function is training-free and rule-based, ensuring it provides a consistent and efficient signal. Below, we provide a detailed description of the reward design.
Lexical Agreement. Inspired by existing works in code generation and AVR (codebleu, ; vrepair, ; vulmaster, ), we adopt BLEU to approximate token similarity between patches. BLEU measures the proportion of overlapping $n$-grams, providing a baseline reward for using correct keywords and identifiers to mitigate token-level hallucination. Formally, the BLEU score for a generated patch $\hat{y}$ against an oracle patch $y$ is defined as:
(2) $\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}\tfrac{\log p_{n}}{N}\right),$
where $p_{n}$ is the $n$-gram precision, and BP is the brevity penalty to discourage excessively short generations.
@@ 3.2.2. Reward
Syntactic Agreement. This aspect of the reward measures how closely the generated patch’s structure matches the oracle’s structure via the AST similarity measure. The AST represents the code’s hierarchical syntax, where each node corresponds to a syntactic construct, and edges encode their hierarchical relationships. Each code snippet is decomposed into a set of subtrees, and similarity is computed as the proportion of matched subtrees. Formally, given S(y^)S($\hat${y}) and $S(y)$ as the sets of subtrees from a generated patch $\hat{y}$ and an oracle patch $y$ respectively, we define AST similarity as:
(3) $\text{Sim}_{\text{AST}}(\hat{y},y)=\frac{|\,S(\hat{y})\cap S(y)\,|}{|\,S(y)\,|}.$
@@
Semantic Agreement. In addition to AST, we leverage the DFG to capture semantic consistency in terms of variable usage and data dependencies. The DFG represents how values propagate through a program: each node corresponds to a variable, and directed edges denote data dependencies. Each code snippet is decomposed into a set of data-dependency tuples in the form $(v_{i},v_{j})$, representing a flow edge from variable $v_{i}$ to $v_{j}$. We then measure similarity as the fraction of matched edges:
(4) $\text{Sim}_{\text{DFG}}(\hat{y},y)=\frac{|\,E(\hat{y})\cap E(y)\,|}{|\,E(y)\,|},$
where E(y^)E($\hat${y}) and $E(y)$ denote the sets of data-flow edges extracted from $\hat{y}$ and $y$, respectively.
@@
Final Reward Computation. We define a reward vector $\mathbf{r}=\left\langle\text{BLEU}(\hat{y},y),\text{Sim}_{\text{AST}}(\hat{y},y),\text{Sim}_{\text{DFG}}(\hat{y},y)\right\rangle$, which components are these agreement scores. We normalize these scores into a single reward ranging from [0, 1] by taking their mean: $Re(\hat{y},y)=\frac{\|\mathbf{r}\|_{1}}{n}$, where $\|\mathbf{r}\|_{1}$ is the L1norm of the vector, representing the sum of the absolute values of its components. This balanced reward function encourages the model to generate patches that are lexically faithful, syntactically consistent, and semantically coherent with the oracle.
3.3. GRPO Optimization
In the RL stage, we optimize the policy model using Group Relative Policy Optimization (GRPO) (shao2024deepseekmath, ) with the reward signal $Re(\hat{y},y)$, which is applied exclusively during training phase with training dataset. In each iteration, we rollout the policy to explore repair spaces, assign semantic-aware rewards to the patches, and update the policy to favor the patches with higher reward.
3.3.1. Overall Process.
Each update iteration of GRPO follows three steps:
  • Rollout. For each vulnerable function $P_{i}$, the current policy $\pi_{\theta}$ generates $M$ candidate outputs in the reasoning–then–edit format $\{(r_{i1},\hat{y}_{i1}),\dots,(r_{iM},\hat{y}_{iM})\}$. These represent diverse repair hypotheses explored by the policy.
  • Reward and Advantage Estimation. Each candidate $\hat{y}_{ij}$ is scored by our rewarder $Re$, yielding $R_{ij}=Re(\hat{y}_{ij},y_{i})$. To account for reward scale differences across inputs, GRPO normalizes scores within each group. Let $\mu_{i}$ and $\sigma_{i}$ be the mean and standard deviation of the $M$ rewards for prompt $P_{i}$, then the normalized advantage is $A_{ij}=\frac{R_{ij}-\mu_{i}}{\sigma_{i}+\epsilon}$, where $\epsilon$ is a small constant used in GRPO implementation for numerical stability. Intuitively, a patch is considered good if its reward is higher than its siblings for the same vulnerable function.
  • Policy Update. The policy is updated to increase the likelihood of candidate patches with higher relative advantage derived from the reward. This is done using importance ratios between new and old policies, clipped to avoid large updates, and weighted by the advantages.
3.3.2. Policy Objective.
@@ 3.3. GRPO Optimization
Formally, let $r_{ij}(\theta)=\frac{\pi_{\theta}(\hat{y}_{ij}\,|\,P_{i})}{\pi_{\theta_{\text{old}}}(\hat{y}_{ij}\,|\,P_{i})}$ be the importance weight, which measures how much more likely the new policy is to generate $\hat{y}_{ij}$ compared to the old one. The GRPO surrogate loss is:
@@
(5) $\mathcal{L}_{\text{GRPO}}(\theta)=\mathbb{E}_{i,j}\Big[\min\big(r_{ij}(\theta)\,A_{ij},\;\operatorname{clip}(r_{ij}(\theta),\,1-\epsilon_{c},\,1+\epsilon_{c})\,A_{ij}\big)\Big]-\beta\,\mathrm{KL}\!\left[\pi_{\theta_{\text{old}}}\,\|\,\pi_{\theta}\right],$
where $\epsilon_{c}>0$ is the clipping threshold that restricts the change of importance ratios, preventing overly aggressive policy updates, and $\beta\geq 0$ is the coefficient of the KL-regularization term that penalizes significant deviations from the previous policy.
3.3.3. Outcome.
Through iterative rollouts, group-relative reward normalization, and clipped updates, GRPO steers the policy toward generating patches that are lexically faithful, syntactically consistent, and semantically coherent with the oracle.
3.4. Curriculum Learning
@@
Our analysis in Section 2.2.3 confirms that coordinating edits across multiple code hunks is a primary challenge for AVR models, causing significant performance degradation. To address this, we adopt a curriculum learning strategy that organizes training from simple to complex repairs. This allows the model to gradually master the skill of performing consistent, multi-hunk edits.
3.4.1. Difficulty Definition.
We use the number of vulnerable hunks as a proxy for repair difficulty, which correlates with the cognitive complexity of the task. We define three stages: easy (1-2 hunks), medium (3-5 hunks), and hard (¿5 hunks). Notably, our “easy” stage intentionally includes two-hunk functions. Preliminary experiments showed that training only on single-hunk fixes caused the model to overfit to overly localized repairs. By including simple multi-hunk cases from the start, we ensure the model begins learning the principles of cross-hunk coordination early in its training.
@@
Our curriculum proceeds through the three stages during the RL phase. The training set is expanded cumulatively at each stage to prevent catastrophic forgetting, and the model’s policy is always initialized from the previous stage’s checkpoint.
3.4.2. Outcome.
This curriculum aligns the training process with the natural hierarchy of repair complexity. By first mastering a foundation of simple, common repair patterns, the model is better equipped to learn the more complex compositions and nuanced logic required for difficult, multi-hunk vulnerabilities, leading to more robust and generalizable performance across the board. Implementation-wise, our curriculum learning is seamlessly integrated with the RL pipeline (Sec. 3.2), as it gradually controls how the training data is presented to the RL process.
4. Experiment Setup
4.1. Implementation Details
We implement SeCuRepair with HuggingFace Transformers (transformers, ) for SFT training, vllm (vllm, ) for inference, and Verl (verl, ) for RL training. Due to excessive training costs for RL, we select one base model, Qwen2.5-7B-Instruct (bai2023qwen, ), as proof-of-concept. Qwen2.5 is commonly used in software engineering tasks (yang2024acecode, ; aggarwal2025nextcoder, ; ke2025niodebugger, ) and relevant vulnerability detection task (weyssow2025r2vul0, ) as the base model. This setting aligns with existing AVR approaches (vrepair, ; favor, ), which fine-tune one base model.
4.2. Dataset
For fair comparison with prior AVR approaches, we adopt the BigVul corpus (fan2020ac, ) as used by VulMaster (vulmaster, ) and FAVOR (favor, ) for the main evaluation. We reuse the deduplicated BigVul split from FAVOR, yielding 9,333 vulnerable functions paired with corresponding patches in C/C++. We construct a repository-level split: all functions from a given repository appear in exactly one of the training, validation, and test sets (no repository overlap). We use an 8:1:1 train/val/test ratio and verify that no repository crosses splits. Following existing AVR approaches (favor, ; vulmaster, ), each vulnerable hunk in the input is bracketed with localization markers <vul_start> and <vul_end>.
To rigorously test how well models generalize to entirely unseen projects, we constructed PrimeVulAVR, a new AVR test dataset derived from PrimeVul (primevul, ), a corpus originally created for vulnerability detection in C/C++. Following the data extraction pipeline from VREPAIR (vrepair, ), we created vulnerable-patched function pairs from PrimeVul’s CVE-linked commits. To ensure a truly external test set, we then filtered out any function pair whose repository was present in our BigVul training set. This process resulted in a clean, out-of-distribution test set containing 1,554 C/C++ function pairs for evaluating cross-repository generalization.
4.3. Baselines and Evaluation Metrics
We evaluate SeCuRepair against the following groups of baselines:
  • Learning-based AVR approaches. We select two state-of-the-art AVR approaches, VulMaster (vulmaster, ) and FAVOR (favor, ). VulMaster applies CWE expert knowledge to guide the repair process and can handle long input sequences. FAVOR augments the input function with CFG and historical patches. We retrain both on our repository-level split using their recommended hyperparameters.
  • Commercial LLMs. We also evaluate against a top-tier commercial model GPT-4o (achiam2023gpt, ) to benchmark against the general-purpose state-of-the-art. We prompt GPT-4o with the same instructional wrapper used for SeCuRepair to ensure a fair, direct comparison of repair capabilities.
  • SeCuRepair base model with SFT. We fine-tune the base model of SeCuRepair, Qwen2.5-7B-Instruct (bai2023qwen, ), on our repository-level BigVul training split as one baseline. We select the best checkpoint based on the performance of the validation set.
On Excluding Agentic Frameworks. We considered including agent-based program repair frameworks as baselines given their recent success. However, the standard agentic loop, which iterates over code localization, repair generation, and execution-based verification (bouzenia2024repairagent, ; tao2024magis, ; antoniades2024swe, ), does not apply to the function-level AVR task. Vulnerabilities are already localized by definition, and providing execution feedback from a potentially vulnerable system is often disallowed for general security applications. Consequently, the agentic loop collapses to a single repair-generation step. This is functionally equivalent to our zero-shot evaluation of GPT-4o. We therefore report the GPT-4o results directly without additional agent scaffolding.
Automatic Evaluation Metrics. We use CodeBLEU as our primary metric following existing works (vulmaster, ; favor, ). Differently, we do not use EM as discussed in our empirical analysis (Section 2.2.2).
5. Experiment Results
5.1. Research Questions
Our experimental evaluation is designed to answer the following research questions:
  • RQ1: How effective is SeCuRepair in repairing vulnerabilities?
  • RQ2: How reasoning-transferred SFT improves repair quality and interpretability?
  • RQ3: How does semantic-aware RL contribute to the performance?
5.2. RQ1. Effectiveness of SeCuRepair
5.2.1. Automatic Evaluation.
Experiment Setting. To evaluate the overall effectiveness of SeCuRepair, we compare it against all baselines on BigVul and our crafted PrimeVulAVR test set (detailed in Section 4.2). For a controlled comparison, all open-source models (our baselines and SeCuRepair) are retrained on our BigVul training split. During evaluation, we use deterministic decoding by setting the temperature to 0 for all models.

### v2

@@ 4 Approach
These principles are realized through a two-stage training pipeline, as shown in Figure 2. In the first stage, we bootstrap the model’s reasoning ability. We employ a commercial teacher LLM to generate high-quality (reasoning, patch) examples with heuristic-based rejection sampling, and then perform SFT on the student model. In the second stage, we refine the model’s patching ability using GRPO algorithm Shao et al. (2024). The model is optimized through a semantic-aware RL optimization, which uses rewards measuring lexical, syntactic, and semantic correctness, and is guided by difficulty-based curriculum training, which gradually increases task complexity to multi-hunk fixes.
@@
Figure 2: Overview of the SeCuRepair training pipeline. (1) Stage 1: Reasoning-enhanced Initialization. A teacher LLM generates (reasoning, patch) candidates, which are filtered via rejection sampling to create a high-quality dataset. The student model is then fine-tuned to learn the reason-then-edit format. (2) Stage 2: Semantics-Guided RL. The SFT model is refined using RL. The policy is rewarded based on a composite score of lexical (BLEU), syntactic (AST), and sematic (CFG & DFG) similarity to the oracle. Stage 2 is organized with a curriculum, progressively training the model on repairs with an increasing number of hunks to master multi-hunk fixes.
4.1 Reasoning-enhanced Initialization
The first stage of our pipeline aims to initialize the model to follow a reason-then-edit paradigm through knowledge distillation.
Knowledge Distillation. To achieve this, the first step is to construct a dataset of (reasoning, patch) pairs by distilling reasoning knowledge from a teacher LLM (i.e., GPT-5 mini). To fully exploit the reasoning capabilities of teacher LLMs, we define a structured reasoning protocol that mirrors the cognitive workflow of human security experts: (1) Vulnerability Classification: characterize the CWE type and security impact; (2) Root Cause Analysis: trace the logical flaw triggering the vulnerability; and (3) Repair Planning: formulate a high-level repairing strategy before patching. Extracted reasoning chains that adhere to this protocol effectively transfer the teacher’s security diagnostic capabilities to the student model.
Rejection Sampling. While the expert workflow guides the reasoning process, it does not guarantee correctness. LLMs often produce plausible-sounding but functionally incorrect patches Dong et al. (2025). Therefore, we adopt a two-step rejection-sampling strategy to denoise reasoning data: 1) we apply syntactic filtering to discard responses that violate the output schema; 2) we apply semantic filtering to retain data only if its patch is similar to the human-reference patch using CodeBLEU as a heuristic proxy.
@@ 4.1 Reasoning-enhanced Initialization
SFT. After rejection sampling, we fine-tune the student model with a standard SFT objective. The resulting model is capable of generating reasoning before patching and serves as the initial policy for the subsequent reinforcement learning stage.
4.2 Semantics-Aware RL
Following SFT stage, SeCuRepair employs RL with semantics-aware rewards, enabling the model to explore the solution space and reward semantic equivalent solutions over rigid textual identity.
@@ 4.2 Semantics-Aware RL
We employ an on-policy RL algorithm GRPO Shao et al. (2024) for this stage. During training, the model generates a (reasoning, patch) pair for a given vulnerability. We then compute a scalar reward $Re$ based on the generated patch against its human reference. This semantics-aware reward moves beyond token overlap to measure the lexical (BLEU), syntactic (Abstract Syntax Tree, AST), and semantic ( DFG and CFG) similarities. It is then used to update the model’s policy, encouraging it to generate patches that achieve higher semantic agreement with human reference.
4.2.1 Reward Design
Our reward assesses the agreement between generated patch and the reference at three granularities:
Lexical Agreement. We calculate BLEU score between the generated and reference patches, which measures the proportion of overlapping $n$-grams. It introduces a token-level reward signal that incentivizes the use of correct keywords and identifiers, thereby suppressing token-level hallucinations. BLEU score for a generated patch $\hat{y}$ against an oracle patch $y$ is defined as: $\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}\tfrac{\log p_{n}}{N}\right),$ where $p_{n}$ is the $n$-gram precision, and BP is the brevity penalty to discourage excessively short generations.
@@ 4.2.1 Reward Design
Syntactic Agreement. AST similarity evaluates the structural alignment between the generated patch and the oracle. Each code snippet is decomposed into a set of subtrees, and similarity is computed as the proportion of matched subtrees. Formally, given S(y^)S($\hat${y}) and $S(y)$ as the sets of subtrees from a generated patch $\hat{y}$ and an oracle patch $y$ respectively, we define AST similarity as: $\text{Sim}_{\text{AST}}(\hat{y},y)=\frac{|\,S(\hat{y})\cap S(y)\,|}{|\,S(y)\,|}$.
@@
Semantic Agreement. We measure semantic consistency from complementary perspectives: CFG (control-flow graph) and DFG (data-flow graph) similarity. The DFG represents how values propagate through a program. Each code snippet is decomposed into a set of data-dependency tuples $(v_{i},v_{j})$. We measure similarity as the fraction of matched edges: $\text{Sim}_{\text{DFG}}(\hat{y},y)=\frac{|\,E(\hat{y})\cap E(y)\,|}{|\,E(y)\,|}$, where E(y^)E($\hat${y}) and $E(y)$ denote the sets of data-flow edges extracted from $\hat{y}$ and $y$, respectively.
Complementarily, the CFG captures the algorithmic structure and execution logic. We extract CFGs using the Joern parser, where nodes represent statements and edges denote control transfers. However, computing precise structural similarity metrics, such as the Graph Edit Distance (GED) with structural matching (i.e., ignoring variable naming differences), is an NP-hard. This computational complexity makes GED inappropriate for online RL, which requires instant reward feedback. To address this efficiency bottleneck, we employ the Weisfeiler-Lehman (WL) Graph Kernel Shervashidze et al. (2011) as a fast, polynomial-time approximation. The WL kernel maps each graph to a high-dimensional feature vector $\phi(G)$ by iteratively aggregating neighborhood labels to capture topological sub-structures (e.g., chains, cycles). We define the CFG similarity as the normalized kernel value: $\text{Sim}_{\text{CFG}}(\hat{y},y)=\frac{k_{\text{WL}}(\hat{y},y)}{\sqrt{k_{\text{WL}}(\hat{y},\hat{y})\cdot k_{\text{WL}}(y,y)}}$.
@@
Final Reward. We first define a syntactic check to ensure patches are structurally valid and can be parsed into AST via Tree-sitter; failed patches receive a zero reward. We then define a reward vector $\mathbf{r}=\langle\text{BLEU},\text{Sim}_{\text{AST}},$ $\text{Sim}_{\text{DFG}},\text{Sim}_{\text{CFG}}\rangle$ for valid fix, where the components represent these agreement scores. We normalize these scores into a single reward ranging from [0, 1] by taking their mean: $Re(\hat{y},y)=\frac{\|\mathbf{r}\|_{1}}{n}$, where $\|\mathbf{r}\|_{1}$ is the L1-norm of the vector, representing the sum of the absolute values of its components. This balanced reward function encourages the model to generate patches that are lexically faithful, syntactically consistent, and semantically coherent with the oracle.
4.3 GRPO Optimization
@@
In the RL stage, we optimize the policy model using Group Relative Policy Optimization (GRPO) Shao et al. (2024) with the reward signal $Re(\hat{y},y)$. Let $r_{ij}(\theta)=\frac{\pi_{\theta}(\hat{y}_{ij}\,|\,P_{i})}{\pi_{\theta_{\text{old}}}(\hat{y}_{ij}\,|\,P_{i})}$ be the importance weight, which measures how much more likely the new policy is to generate $\hat{y}_{ij}$ compared to the old one. The GRPO surrogate loss is:
@@
$\mathcal{L}_{\text{GRPO}}(\theta)={}$ $\mathbb{E}_{i,j}\Big[\min\big(r_{ij}(\theta)\,A_{ij},\;\operatorname{clip}(\epsilon_{c})\,A_{ij}\big)\Big]$
${}-\beta\,\mathrm{KL}\!\left[\pi_{\theta_{\text{old}}}\,\|\,\pi_{\theta}\right]$
4.4 Curriculum Learning
@@
Our analysis in Section 3.3 confirms that coordinating edits across multiple code hunks is a primary challenge for AVR models. To address this, we adopt a curriculum learning strategy that organizes training from simple to complex repairs. This approach decouples the learning process by progressively increasing task complexity, allowing the model to master local security edits before tackling global syntactic synchronization.
@@
We use the number of code hunks in human reference as a proxy for repair difficulty. We define three stages: easy (1-2 hunks), medium (3-5 hunks), and hard (>5 hunks). Preliminary experiments showed that training only on single-hunk fixes caused the model to overfit to overly localized repairs; therefore, our easy stage includes two-hunk functions. Our curriculum proceeds through the three stages during the RL phase. The training set is expanded cumulatively at each epoch to prevent catastrophic forgetting, and the model’s policy is always initialized from the previous stage’s checkpoint.
5 Experimental Setup
5.1 Baselines and Evaluation Metrics
We benchmark SeCuRepair against specialized AVR approaches (VulMaster Zhou et al. (2024), FAVOR Dong et al. (2025)) retrained on our repository-level split, a commercial SOTA model (GPT-4o Achiam et al. (2023)) prompted with identical instructions, and our own SFT-only base model (Qwen2.5-7B) to isolate the contributions of the RL stage. More details are in Appendix G.
5.2 Datasets
To strictly prevent data leakage and evaluate cross-project generalization, we enforce a repository-level split where no repository overlaps between training, validation, and test sets. Following selected baselines, we utilize BigVul Fan et al. (2020). In addition, we construct a new out-of-distribution test set PrimeVulAVR derived from the vulnerability detection dataset PrimeVul Ding et al. (2024). We rigorously filtered out any repositories present in the BigVul training set, resulting in 1,554 unique function pairs. Detailed construction pipelines and statistics are provided in Appendix E.
5.3 Implementation Details
We implement SeCuRepair based on Qwen2.5-7B-Instruct Bai et al. (2023) using HuggingFace 26 and Verl 27. For SFT, we perform full-parameter fine-tuning for 3 epochs with a learning rate of $3\times 10^{-5}$ and a context length of 4,096, computing loss exclusively on LLM response. For RL, we employ the GRPO algorithm initialized from the best SFT checkpoint. We set the global batch size to 1,024 with $M=8$ rollouts per prompt and train for up to 20 epochs with an actor learning rate of $1\times 10^{-6}$. Full details are available at Appendix F. We use CodeBLEU as our primary metric following existing works Zhou et al. (2024); Dong et al. (2025). We exclude the Exact Match (EM) metric from our main evaluation. Under the rigorous repository-level split, all models achieve negligible scores ($5\%$), rendering the metric indistinguishable (more discussion of EM is in Appendix H).
6 Results
We evaluate the effectiveness of SeCuRepair in repairing vulnerabilities under both automatic and human evaluation. The automatic evaluation results are summarized in Table 1, with the human study results in Table 2. We present the ablation performance of reasoning in Table 3 and Figure 4, semantic-aware RL in Figure 3, and curriculum training in Figure 5.
1) SeCuRepair performs better than SOTA baselines. Table 1 shows that SeCuRepair outperforms all baselines on both datasets by a significant margin in terms of CodeBLEU. Specifically, SeCuRepair surpasses the best-performing baseline on BigVul, VulMaster, by 37.21%; and surpasses the best-performing baseline on PrimeVulAVR, GPT-4o, by 33.58%. This result demonstrates the superior ability of SeCuRepair to repair vulnerability and generalize to unseen repositories. A Wilcoxon signed-rank test Woolson (2007) confirms that all of SeCuRepair’s performance gains over the baselines are statistically significant (p < 0.001).
Notablly, SeCuRepair outperforms the monolithic SFT baseline (Qwen2.5-7B+SFT) with CodeBLEU gains of 21.98% on BigVul and 20.64% on PrimeVulAVR, confirming the superiority of our training framework over standard SFT.
2) Human evaluation validates the superior performance of SeCuRepair. To complement automatic metrics, we conducted a blind human evaluation comparing SeCuRepair against VulMaster and GPT-4o. We recruited four experts with experience in software security and C++ to evaluate 309 statistically sampled instances (95% confidence level). Participants rated patch semantic similarity on a 5-point Likert scale relative to the human reference (full detail is available at Appendix I). Aligning with the goal of AI-assisted repair Takerngsaksiri et al. (2025), we categorize scores $\geq 3$ as “workable drafts,” indicating patches that capture sufficient logic to serve as valid starting points for developers.